# Model Stock: All we need is just a few fine-tuned models

### v1

@@ 2.1 Geometric Properties Between Weights
Interestingly, these consistencies in angle and norm are observed 1) across diverse setups and 2) both during and after training. Fig. 2 shows this consistency over 50 fine-tuned CLIP ViT-B/32333https://github.com/openai/CLIP [24] models. It illustrates that the layer-wise norm and angle of these models exhibit almost constant values with extremely minimal error. While this figure depicts a specific model (i.e., CLIP ViT-B/32), we establish that such regularity is not confined to a single model or setting but is consistent across various CLIP fine-tuning scenarios. We conjecture this holds irrespective of networks (ViT [5], Hybrid-ViT [5], ResNet [9], ConvNext [21]), optimizers (SGD, AdamW [22]), data augmentations (RRC [27], RandAug [4]), datasets (CIFAR [14], ImageNet [26]), or initialization of the classifier (zero-shot, LP-FT [15]). Remarkably, this regularity also holds for fine-tuned weights at each step during training as well as after training. A comprehensive analysis supporting these findings is presented in the Appendix.
Based on the observation, we presume the distribution of the fine-tuned weights. The center of the fine-tuned weights is defined as $\bm{\mu}$$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\mathbf{w}_{i}$. We then deduce the following properties among fine-tuned weights: (i) $\|\mathbf{w}_{i}-\bm{\mu}\|=\text{constant}$, indicating a thin shell distribution; (ii) $(\mathbf{w}_{0}-\bm{\mu})\perp(\mathbf{w}_{i}-\bm{\mu})$; and (iii) $(\mathbf{w}_{i}-\bm{\mu})\perp(\mathbf{w}_{j}-\bm{\mu})$ for all $i,j\in[1,N]$. These properties are depicted in Fig. 5 for better understanding. The detailed proof is in the Appendix.
@@ 2.2 Center of Weights and Performance
We present an additional observation regarding the test error landscape, which relates to the performance around the center of weights. Fig. 3 depicts the test error landscape on the ImageNet test dataset within a two-dimensional plane. This plane includes a pre-trained model’s weight ($\mathbf{w}_{0}$), a single fine-tuned model ($\mathbf{w}_{1}$), and the pseudo-center ($\mathbf{w}_{\text{avr}}^{(50)}$). This landscape reveals that a fine-tuned model typically occupies the boundary of test error regions. On the other hand, centered near pseudo-center ($\mathbf{w}_{\text{avr}}^{(50)}$), the test error is the lowest and gets higher as the weights get far from the center. Interestingly, along the line segment $\overline{\mathbf{w}_{0}\mathbf{w}_{1}}$ (i.e., a WiSE-FT curve [33]), the fine-tuned weight $\mathbf{w}_{1}$ is neither the point closest to the pseudo-center nor the one with the lowest test error. We will connect this observation in §3 and §5 to find $\mathbf{w}_{H}$, the weight on the line closest to the center.
@@ 3 Method
Figure 6: Schematic concept of our Model Stock. We present two scenarios with a small angle (left) and a large angle (right). Given a pre-trained weight ($\mathbf{w}_{0}$) and two fine-tuned weights ($\mathbf{w}_{1}$, and $\mathbf{w}_{2}$), we visualize a gray triangle representing the span of three weights. We consider this triangle area our search space spanned by three weights. We aim to find the best weight point on the triangle nearest the ideal center $\bm{\mu}$. We find that the perpendicular foot $\mathbf{w}_{H}$ from point $\bm{\mu}$ to the plane is the nearest point, which can be specified solely using the angle between the fine-tuned models, even without knowing the exact position of the center $\bm{\mu}$. We utilize $\mathbf{w}_{H}$ as the merged weight of Model Stock. Intuitively, when the angle $\theta$ is large (i.e. two weights are diverse), as in right figure, $\mathbf{w}_{H}$ will rely more on the pre-trained weight($\mathbf{w}_{0}$), vice versa.
@@
For more detailed proof, please refer to the Appendix. Note that the interpolation ratio $t=\frac{2\cos\theta}{1+\cos\theta}$ is solely determined by the angle $\theta$ between two fine-tuned models. Crucially, unlike previous methods, determining $t$ does not require extra training [32, 19, 28, 29] or heuristic hyper-parameter settings [6, 32], thereby simplifying the process and enhancing its accessibility and efficiency.
It is important to note that as the $\theta$ decreases, the pre-trained model is less utilized for merging, as shown in Fig 6 (left). Coupled with the observation in Fig. 2, this indicates that bias layers rely less on pre-trained models and focus more on fine-tuned models, whereas weight layers depend more on pre-trained models. This observation extends the findings of previous works such as BitFit [34] and LP-FT [15]. In the case of BitFit and LP (i.e., the first step of LP-FT), bias and classifier layers fully utilize fine-tuning, while other weight layers (attention and MLP) rely on pre-trained models.
@@
More detailed proof is in the Appendix. Similar to the case of two fine-tuned models, the interpolation ratio $t$ depends solely on the angle $\theta$ between the pre-trained model and the $N$ fine-tuned models.
Periodic merging. To move one step forward with our method, we propose periodic merging, which is performed between the fine-tuned models and the pre-trained model during training. As the geometric properties of weights are also applicable to weights during training (refer to the Appendix for more details), we fully utilize this phenomenon here.
@@
We strategically merge weights at the end of every epoch, allowing for the fine-tuning process to be parallelized with merging. The interpolation ratio for merging is determined by the angle $\theta$ between the pre-trained model and the fine-tuned models at the current epoch. Fig. 7 visualizes this periodic merging process. We argue that employing the periodic merging can approximate the center of weights more accurately. In §4.3, we empirically show that the periodic merging yields superior performance and achieves a closer distance to the center.
@@ 4 Experiment
We present the key experimental results in this section. We first provide our experimental setups in §4.1. Then, we present the main results in §4.2, and ablation studies in §4.3. More detailed experimental setups, analysis, and further results are in the Appendix. We will release our codes and weights publicly.
@@ 4.1 Experimental Setup
We fine-tune models on the ImageNet-1K [26] training dataset. We report ImageNet-1K [26] top-1 accuracy for evaluating in-distribution (ID) performance. For distribution shift scenarios, we consider five out-of-distribution (OOD) benchmarks including ImageNet-V2 [25], ImageNet-R [10], ImageNet-Sketch [31], ImageNet-A [11], and ObjectNet [1]. Since previous methods, except Model Soups [32], have not been evaluated on ObjectNet, we omit the ObjectNet results when comparing against them.
@@
We initialize the classifier weights (e.g., 1000 classes for ImageNet) using the text encoder of CLIP and text prompts following Model Soup [32]. We use AdamW [22] with batch size 512 and weight decay 0.1 for all the experiments, including vanilla fine-tuning, Model Soup reproduction, and Model Stock  [^4] We reduce the batch size to 64 on CLIP ViT-L/14 due to memory limitation.. We employ two training setups for comparisons: (1) training 10 epochs with minimal data augmentation, following Model Soup’s zero-shot initialization setup, and (2) training 16 epochs with strong data augmentation, following one of Model Soup’s LP initialization setup [32], denoting with . These setups enable a balanced comparison of Model Stock against zero-shot and LP-initialized Model Soups. Detailed hyper-parameters are in the Appendix.
@@
4.2 Main Results
4.2.1 CLIP ViT-B/32.
Table 2 shows the results of Model Stock on the pretrained CLIP ViT-B/32 model by comparing with Model Soups. ‘Avg. shifts’ denotes the average accuracy of the five OOD benchmark scores. Our Model Stock and Model Stock show competitive performance with Model Soups. Furthermore, Model Stock achieves a state-of-the-art performance on ImageNet with 81.19% top-1 accuracy. As described in Fig. 1, Model Stock with WiSE-FT [33] enjoys superior ID-OOD performance curve compared to Model Soup and its WiSE-FT curves. Note that Model Soups require dozens of fine-tuned models (e.g., zero-shot and LP-init Model Soups use 48 and 71 models, respectively), highlighting the effectiveness of Model Stock along with efficiency utilizing only two models. We provide further comparison results with WiSE-FT [33] curves on LP-init Model Soups in the Appendix.
@@ 4.2 Main Results
Table LABEL:exp:vit_b_16 presents a comprehensive comparison of different fine-tuning methods applied to CLIP ViT-B/16. Previous works [15, 28, 29] lack ObjectNet [1] results; therefore, we omit ObjectNet and report the other four OOD benchmarks. Complete results with ObjectNet and ImageNet-Real [2] are in the Appendix. The results show Model Stock exhibits exceptional performance on ImageNet accuracy, e.g., Model Stock achieves 85.2% top-1 accuracy on ImageNet, which is a state-of-the-art level. Model Stock also shows robust performance across diverse distribution shift scenarios. We emphasize the simplicity of Model Stock devoiding additional fine-tuning tricks, such as knowledge distillation [23] or additional optimization [29]. This highlights the general applicability of Model Stock.
@@ 4.3 Ablation studies and analysis of Model Stock
4.3.1 Experiments on the number of fine-tuned model $N$.
4.3.2 Study on the merging period of Model Stock.
4.3.3 The post-training merging strategy of Model Stock.
5 Discussion — Rethinking Pivotal Prior Studies
In this section, we extend our findings to reinterpret the underlying mechanics in prior studies [33, 32]. Building upon our knowledge above, we offer insights into two milestone methods, WiSE-FT [33] and Model Soups [32], through a consistent rationale to illuminate their effectiveness.
5.0.1 WiSE-FT
[33] is a state-of-the-art robust fine-tuning method for CLIP-based models. It demonstrates that linearly combining weights of the pre-trained and fine-tuned models acheive significant accuracy gain on distribution shifts and ImageNet. We argue that the WiSE-FT model’s superiority over a fine-tuned model can be interpreted by its weights being closer to the center of the corresponding weight distribution. Fig. 3 already showed fine-tuned models typically lie on the periphery of flat minima. Given that the angle $\angle\textbf{w}_{0}\textbf{w}_{\text{avr}}^{(50)}\textbf{w}_{1}$ is nearly a right angle, along the line $\overline{\textbf{w}_{0}\textbf{w}_{1}}$, multiple weight points are closer to the center than a single fine-tuned model, thereby enhancing performance. Note that $\textbf{w}_{H}$ is the closest to the center among the line $\overline{\textbf{w}_{0}\textbf{w}_{1}}$. More discussions on performance boosts observed in distribution shifts are provided in the Appendix.
5.0.2 Model Soup
[32] merges various fine-tuned models’ weights trained from varied hyper-parameters. It has been credited with delivering enhanced performance across ImageNet and distribution shifts. Here, we interpret the performance improvements of Model Soup as the result of the proximity to the center of weight distribution. Consider two weight vectors, $\mathbf{w}_{A}$ and $\mathbf{w}_{B}$, fine-tuned with different hyper-parameters and following Gaussian distribution $\mathcal{N}(\bm{\mu_{A}},\Sigma_{A})$ and $\mathcal{N}(\bm{\mu_{B}},\Sigma_{B})$ respectively. The interpolated weight vector $\mathbf{w}_{AB}=t\cdot\mathbf{w}_{A}+(1-t)\cdot\mathbf{w}_{B}$ also follows a Gaussian distribution $\mathcal{N}(\bm{\mu_{AB}},\Sigma_{AB})$. The expected squared distance from the interpolated weight vector to its mean $\bm{\mu_{AB}}$ is minimized to $\frac{\text{trace}(\Sigma_{A})\text{trace}(\Sigma_{B})}{\text{trace}(\Sigma_{A})+\text{trace}(\Sigma_{B})}$ when $t$ is chosen to $\frac{\text{trace}(\Sigma_{B})}{\text{trace}(\Sigma_{A})+\text{trace}(\Sigma_{B})}$, indicating the reduction of variance through weight interpolation (i.e., the distance between $\mathbf{w}_{AB}$ and $\bm{\mu_{AB}}$ might be closer than each weight’s distance). For example, if $\text{trace}(\Sigma_{B})$ is equal to $\text{trace}(\Sigma_{A})$, this minimum squared distance is exactly half of the sum of the individual traces when $t=0.5$. This insight suggests that the performance gains realized by Model Soup could be due to reduced variance resulting from merging numerous weights.
We conduct a toy experiment that we assess the ImageNet performance of interpolated weights between pairs of fine-tuned models within Greedy Model Soup [^5] We opt for Greedy Model Soup to demonstrate that even the ensemble of the best-performing models does not benefit from the ensembling impact. [32] and compare it to interpolations between their central weights, calculated as the average of 20 differently seeded models. Fig. 8 shows that, unlike interpolations between individual models, using the centers does not improve performance. Our results indicate that proximity to the center of the weight distribution may play a more critical role in enhancing performance than the benefits of hyper-parameter diversity in weight ensemble methods.
Figure 8: Ensembling impact disappears when interpolating between two averaged weights. We plot the ImageNet performance of interpolated weights between two selected fine-tuned models in Model Soup [32] models (left) and between two averaged weights (center) of two fine-tuned model groups (right).
@@ 6 Conclusion
Additionally, our findings suggested further knowledge and applicability to the models near flat minima and offered new perspectives on methods like WiSE-FT and Model Soup. As the pretraining-finetuning paradigm gains more prominence, our insights will provide a foundation for better understanding and optimizing the fine-tuning process in both academia and industry.

### v2

@@ 2.1 Geometric Properties Between Weights
Interestingly, these consistencies in angle and norm are observed 1) across diverse setups and 2) both during and after training. Fig. 2 shows this consistency over 50 fine-tuned CLIP ViT-B/32333https://github.com/openai/CLIP [30] models. It illustrates that the layer-wise norm and angle of these models exhibit almost constant values with extremely minimal error. While this figure depicts a specific model (i.e., CLIP ViT-B/32), we establish that such regularity is not confined to a single model or setting but is consistent across various CLIP fine-tuning scenarios. We conjecture this holds irrespective of networks (ViT [5], Hybrid-ViT [5], ResNet [9], ConvNext [24]), optimizers (SGD, AdamW [25]), data augmentations (RRC [35], RandAug [4]), datasets (CIFAR [17], ImageNet [34]), or initialization of the classifier (zero-shot, LP-FT [18]). Remarkably, this regularity also holds for fine-tuned weights at each step during training as well as after training. A comprehensive analysis supporting these findings is presented in the Appendix 0.A.
Based on the observation, we presume the distribution of the fine-tuned weights. The center of the fine-tuned weights is defined as $\bm{\mu}$$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\mathbf{w}_{i}$. We then deduce the following properties among fine-tuned weights: (i) $\|\mathbf{w}_{i}-\bm{\mu}\|=\text{constant}$, indicating a thin shell distribution; (ii) $(\mathbf{w}_{0}-\bm{\mu})\perp(\mathbf{w}_{i}-\bm{\mu})$; and (iii) $(\mathbf{w}_{i}-\bm{\mu})\perp(\mathbf{w}_{j}-\bm{\mu})$ for all $i,j\in[1,N]$. These properties are depicted in Fig. 5 for better understanding. The detailed proof is in the Appendix 0.B.
@@ 2.2 Center of Weights and Performance
We present an additional observation regarding the test error landscape, which relates to the performance around the center of weights. Fig. 3 depicts the test error landscape on the ImageNet test dataset within a two-dimensional plane. This plane includes a pre-trained model’s weight ($\mathbf{w}_{0}$), a single fine-tuned model ($\mathbf{w}_{1}$), and the pseudo-center ($\mathbf{w}_{\text{avr}}^{(50)}$). This landscape reveals that a fine-tuned model typically occupies the boundary of test error regions. On the other hand, centered near pseudo-center ($\mathbf{w}_{\text{avr}}^{(50)}$), the test error is the lowest and gets higher as the weights get far from the center. Interestingly, along the line segment $\overline{\mathbf{w}_{0}\mathbf{w}_{1}}$ (i.e., a WiSE-FT curve [41]), the fine-tuned weight $\mathbf{w}_{1}$ is neither the point closest to the pseudo-center nor the one with the lowest test error. We will connect this observation in §3 to find $\mathbf{w}_{H}$, the weight on the line closest to the center.
@@ 3 Method
Figure 6: Schematic concept of our Model Stock. We present two scenarios with a small angle (left) and a large angle (right). Given a pre-trained weight ($\mathbf{w}_{0}$) and two fine-tuned weights ($\mathbf{w}_{1}$, and $\mathbf{w}_{2}$), we visualize a gray triangle representing the span of three weights. We consider this triangle area our search space spanned by three weights. We aim to find the best weight point on the triangle nearest the ideal center $\bm{\mu}$. We find that the perpendicular foot $\mathbf{w}_{H}$ from point $\bm{\mu}$ to the plane is the nearest point, which can be specified solely using the angle between the fine-tuned models, even without knowing the exact position of the center $\bm{\mu}$. We utilize $\mathbf{w}_{H}$ as the merged weight of Model Stock. Intuitively, when the angle $\theta$ is large (i.e., two weights are diverse), as in the right figure, $\mathbf{w}_{H}$ will rely more on the pre-trained weight($\mathbf{w}_{0}$), vice versa.
@@
For more detailed proof, please refer to the Appendix 0.D. Note that the interpolation ratio $t=\frac{2\cos\theta}{1+\cos\theta}$ is solely determined by the angle $\theta$ between two fine-tuned models. Crucially, unlike previous methods, determining $t$ does not require extra training [40, 22, 36, 37] or heuristic hyper-parameter settings [6, 40], thereby simplifying the process and enhancing its accessibility and efficiency.
As the $\theta$ decreases, the pre-trained model is less utilized for merging, as shown in Fig 6 (left). Coupled with the observation in Fig. 2, this indicates that bias layers rely less on pre-trained models and focus more on fine-tuned models, whereas weight layers depend more on pre-trained models. This observation extends the findings of previous works such as BitFit [43] and LP-FT [18]. In the case of BitFit and LP (i.e., the first step of LP-FT), bias and classifier layers fully utilize fine-tuning, while other weight layers (attention and MLP) rely on pre-trained models. We present an additional analysis in the Appendix 0.F.
@@
Detailed proof is in the Appendix 0.D. Similar to the case of two fine-tuned models, the interpolation ratio $t$ depends solely on the angle $\theta$ between the pre-trained model and the $N$ fine-tuned models.
@@
Periodic merging. To move one step forward with our method, we propose periodic merging, which is performed between the fine-tuned models and the pre-trained model during training. As the geometric properties of weights are also applicable to weights during training (refer to Appendix 0.A.2 for more details), we fully utilize this phenomenon here.
We strategically merge weights at the end of every epoch, allowing for the fine-tuning process to be parallelized with merging. The interpolation ratio for merging is determined by the angle $\theta$ between the pre-trained model and the fine-tuned models at the current epoch. Fig. 7 visualizes this periodic merging process. We argue that employing a periodic merging can approximate the center of weights more accurately. In §4.3, we empirically show that the periodic merging yields superior performance and achieves a closer distance to the center.
@@ 4 Experiment
We present the key experimental results in this section. We first provide our experimental setups in §4.1. Then, we present the main results in §4.2, and ablation studies in §4.3. More detailed experimental setups, analysis, and further results are in the Appendix 0.G. We will release our codes and weights publicly.
@@ 4.1 Experimental Setup
We fine-tune models on the ImageNet-1K [34] training dataset. We report the ImageNet-1K [34] top-1 accuracy for evaluating in-distribution (ID) performance. For distribution shift scenarios, we consider five out-of-distribution (OOD) benchmarks including ImageNet-V2 [33], ImageNet-R [10], ImageNet-Sketch [39], ImageNet-A [11], and ObjectNet [1]. Since previous methods, except Model Soups [40], have not been evaluated on ObjectNet, we omit the ObjectNet results when comparing against them.
@@
We initialize the classifier weights (e.g., 1000 classes for ImageNet) using the text encoder of CLIP and text prompts following Model Soup [40]. We use AdamW [25] with batch size 512 and weight decay 0.1 for all the experiments, including vanilla fine-tuning, Model Soup reproduction, and Model Stock  [^4] We reduce the batch size to 64 on CLIP ViT-L/14 due to memory limitation.. We employ two training setups for comparisons: (1) training 10 epochs with minimal data augmentation, following Model Soup’s zero-shot initialization setup, and (2) training 16 epochs with strong data augmentation, following Model Soup’s LP initialization setup [40], denoted with . These setups enable a balanced comparison of Model Stock against zero-shot and LP-initialized Model Soups. Detailed hyper-parameters are in the Appendix 0.G.
@@
4.2 Main Results
4.2.1 CLIP ViT-B/32.
Table 2 shows the results of Model Stock on the pre-trained CLIP ViT-B/32 model by comparing it with Model Soups. ‘Avg. shifts’ denotes the average accuracy of the five OOD benchmark scores. Our Model Stock and Model Stock show competitive performance with Model Soups. Furthermore, Model Stock achieves state-of-the-art performance on ImageNet with 81.19% top-1 accuracy. As described in Fig. 1, Model Stock with WiSE-FT [41] enjoys a superior ID-OOD performance curve compared to Model Soup and its WiSE-FT curves. Note that Model Soups require dozens of fine-tuned models (e.g., zero-shot and LP-init Model Soups use 48 and 71 models, respectively), highlighting the effectiveness of Model Stock along with efficiency utilizing only two models. We provide further comparison results with WiSE-FT [41] curves on LP-init Model Soups in the Appendix 0.H.
@@ 4.2 Main Results
Table LABEL:exp:vit_b_16 presents a comprehensive comparison of different fine-tuning methods applied to CLIP ViT-B/16. Previous works [18, 36, 37] lack ObjectNet [1] results; therefore, we omit ObjectNet and report the other four OOD benchmarks. Complete results with ObjectNet and ImageNet-Real [2] are in the Appendix 0.H.2. The results show Model Stock exhibits exceptional performance on the ImageNet accuracy, e.g., Model Stock achieves 85.2% top-1 accuracy on ImageNet, which is a state-of-the-art level. Model Stock also shows robust performance across diverse distribution shift scenarios.
@@ 4.3 Ablation studies and analysis of Model Stock
4.3.1 Experiments on the number of fine-tuned models $N$.
4.3.2 Study on the merging period of Model Stock.
4.3.3 The post-training merging strategy of Model Stock.
@@
5 Related Work
We discuss related works and highlight how our method differs and contributes to the existing works.
5.0.1 Model Soups [40]
is a straightforward weight averaging method that merges weights from various fine-tuned models trained with different hyper-parameters. It demonstrates improved in-distribution (ID) and out-of-distribution (OOD) performance. While effective, model soups typically require a large number of fine-tuned models. Our method aims to achieve similar or superior performance improvements more efficiently, utilizing significantly fewer fine-tuning costs. We provide further discussion about Model Soups with our new interpretation in Appendix 0.E.
5.0.2 Robust Fine-tuning.
When fine-tuning generalist models like CLIP [30], we often observe the fine-tuned models lose the generalization ability of the original ones, with decreased OOD performance. To address this issue, several robust fine-tuning approaches have been proposed. LP-FT [18] attempts to preserve pre-trained weights by initially training only a linear probing layer. WiSE-FT [41] improves OOD performance through linear interpolation between fine-tuned and pre-trained weights. While our method shares similarities with WiSE-FT in using pre-trained weights, our method determines interpolation ratios layer-wise based on geometric properties. Focusing on OOD performance, methods suggesting improved training objectives [27, 29, 28, 7, 36, 37] have been proposed. Our approach differs from these methods as we do not propose a new fine-tuning loss. Instead, we perform two fine-tunings and achieve robust performance by merging them.
5.0.3 Weight Center and Flat minima.
Recent machine learning research has extensively explored the significance of finding flat minima for improved generalization [26, 15, 21, 3]. Keskar et al. [16] and Hochreiter & Schmidhuber [12] demonstrated that sharp optima by large batch SGD have steep, harmful directions, while broader optima enhance generalization. Stochastic Weight Averaging (SWA) [15] targets the center of flat minima, enhancing robustness against shifts in the loss landscape between training and test datasets. SWAG [26] builds on SWA by incorporating Bayesian model averaging with a Gaussian posterior to further boost performance. SWAD [3] found that the generalization gap between flat and sharp minima is more pronounced in OOD scenarios than in ID ones. Our method theoretically extends these approaches by efficiently identifying the center of flat minima with novel geometric properties, leading to significant improvements in both ID and OOD performances.
5.0.4 Model Weight Merging.
Recent research has explored merging models fine-tuned on various tasks. Methods such as Task Arithmetic [14] and TIES [42] have been proposed. They are also based on the difference between fine-tuned and pre-trained weights (often referred to as the “task vector”). However, our method distinguishes itself through geometric analysis for weight merging. In the domain of Large Language Models (LLMs), merging techniques like WARM [31] and WARP [32] have emerged. Our method has the potential for extension to these areas, offering new avenues for future research.
@@ 6 Conclusion
Additionally, our findings suggested further knowledge and applicability to the models near flat minima and will offer new insights on model weight merging methods. As the pretraining-finetuning paradigm gains more prominence, our insights will provide a foundation for better understanding and optimizing the fine-tuning process in both academia and industry.
@@
7 Acknowledgment
We thank the researchers at NAVER AI Lab for their valuable comments. This work was supervised by Sangdoo Yun and Dongyoon Han. Dong-Hwan Jang is currently at Samsung Advanced Institute of Technology (SAIT).